The mean number of births per minute in a country in a recent year was about nine. Find the probability that the

number of births in any given minute is (a) exactly six, (b) at least six, and (c) more than six.


1
Expert's answer
2022-05-24T17:47:27-0400

Let "X=" the number of births : "X\sim Po(\lambda)."

Given "\lambda=9."

(a)


"P(X=6)=\dfrac{e^{-9}(9)^6}{6!}=0.09109"

(b)


"P(X\ge6)=1-P(X=0)-P(X=1)"

"-P(X=2)-P(X=3)"

"-P(X=4)-P(X=5)"

"=1-\dfrac{e^{-9}(9)^0}{0!}-\dfrac{e^{-9}(9)^1}{1!}"

"-\dfrac{e^{-9}(9)^2}{2!}-\dfrac{e^{-9}(9)^3}{3!}"

"-\dfrac{e^{-9}(9)^4}{4!}-\dfrac{e^{-9}(9)^5}{5!}=0.88431"

(c)


"P(X>6)=P(X\ge 6)-P(X=6)"

"=0.79322"